OpenStudy (anonymous):

Solve for x logarithms: log 9 ^5√27=x 3

OpenStudy (anonymous):

Clever way to write a subscript. :)

OpenStudy (anonymous):

Have you learned the log rules? If you are multiplying within a log, it's the same as adding the logs of the two.

OpenStudy (anonymous):

30sqrt(3) ?

OpenStudy (triciaal):

is that 9^5(sqrt^27) in the question?

OpenStudy (anonymous):

the log is included everything under the words

OpenStudy (anonymous):

It's log base 3 of (9^5 x sqrt27) I think.

OpenStudy (anonymous):

Use the multiplication/addition log rule! Then it's easy from there. Do you know of it?

OpenStudy (anonymous):

no the 5 is over the square root sign like ^5√

OpenStudy (anonymous):

no what is it

OpenStudy (anonymous):

It's the fifth root?

OpenStudy (anonymous):

Comme ca?

OpenStudy (anonymous):

yes thats it

OpenStudy (anonymous):

Then you apply the log rule I sent an example of before.

OpenStudy (anonymous):

Tell me what you get after doing that and if you're still unable to solve it then I'll continue.

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

Ah, the image is way complicated. Wait a moment.

OpenStudy (agent0smith):

simplify the 9*(27)^1/5 unfortunately doesn't show well w/o equation editor: 9*(27)^1/5 = 3^2*(3^3)^1/5

OpenStudy (agent0smith):

= 3^(2 + 3/5) = 3^(13/5)

OpenStudy (anonymous):

It says solve for 'X'. So I did log base 3 9+1/5 log base 3 27=x then power rule 2+1/5(3)=X 2.6=x Am I on the right track ? KinzaN your thoughts?

OpenStudy (anonymous):

Ah, you are. There was no need to bring the 1/5 down though. :)

OpenStudy (anonymous):

27^(1/5) can be changed to 3 to some exponent.

OpenStudy (agent0smith):

13/5 = 2.6, so yes.

OpenStudy (triciaal):

is ( 5sqrt27) the power to which 9 is raised? 9 = 3^2 and 27 =3^3 from log rule 3^x = 9 ^5√27

OpenStudy (anonymous):

You can check that link triciaal

OpenStudy (triciaal):

thanks now i can see the problem

OpenStudy (anonymous):

wio?

OpenStudy (anonymous):

Your answer has already been given, if that's what you're looking for. :)

OpenStudy (triciaal):

I don't know what wio means once i was able to see the problem I saw the correct answer was already given

OpenStudy (anonymous):

Still haven't figured it out? =(